### v2

# Evaluating the Impact of Inter-cluster Communications in Edge Computing

Abstract
Distributed applications based on micro-services in edge computing are becoming increasingly popular due to the rapid evolution of mobile networks. While Kubernetes is the default framework when it comes to orchestrating and managing micro-service-based applications in mobile networks, the requirement to run applications between multiple sites at cloud and edge poses new challenges. Since Kubernetes does not natively provide tools to abstract inter-cluster communications at the application level, inter-cluster communication in edge computing is becoming increasingly critical to the application performance. In this paper, we evaluate for the first time the impact of inter-cluster communication on edge computing performance by using three prominent, open source inter-cluster communication projects and tools, i.e., Submariner, ClusterLink and Skupper. We develop a fully open-source testbed that integrates these tools in a modular fashion, and experimentally benchmark sample applications, including the ML class of applications, on their performance running in the multi-cluster edge computing system under varying networking conditions. We experimentally analyze two classes of envisioned mobile applications, i.e., a) industrial automation, b) vehicle decision drive assist. Our results show that ClusterLink performs best out of the three tools in scenarios with increased payloads, regardless of the underlying networking conditions or transmission direction between clusters. It is closely followed by Skupper, unless request and reply both transport significant amounts of data. Finally, when requesting smaller amounts of data from a service, Submariner slightly outperforms Skupper and ClusterLink regardless of the inter-node networking conditions.
Index Terms:
Inter-cluster, Multi-cluster, Kubernetes, Benchmarking, QoS
I Introduction
With the rapid evolution of mobile networks towards 5G/6G, a myriad of new mobile applications are emerging as ever more distributed in nature. Together with cloud computing in the so-called compute continuum, edge computing is becoming the key solution in mobile networks today to providing flexible, low-latency and distributed computation. To deploy highly distributed applications, container orchestration platforms play a critical role, with Kubernetes not only being the most commonly used solution in the cloud, but also increasingly utilized at the edge. Kubernetes organizes multiple nodes into logical clusters, sharing a common control plane and working in union to deploy and run applications. On the other hand, some specialized environments like mobile or Multi-Access Edge Computing (MEC) might introduce volatile nodes that can leave and join or rejoin clusters, creating coordination and service migration overhead. In edge environments, the restriction of cluster size is often necessary to this end, which in turn limits the application size and complexity that can be deployed per cluster. To overcome this issue, service components have to be spread across multiple clusters while still allowing for communication between them.
The inter-cluster communication in the edge context is in fact in its infancy. In this area, research to date majorly focuses on either the placement of workloads across the infrastructure[1], orchestration of multiple clusters[2] or the methods for cluster creation based on various properties, such as node proximity[3]. However, research on how these components can horizontally scale by actually establishing connectivity between them and how this impacts application performance is lacking. Static IP connectivity, while feasible for cloud environments, would require a significant maintenance overhead in dynamic environments like MEC. In fact, such an approach is not suited for large-scale and evolving deployments, therefore necessitating flexible connectivity creation based on service name resolution across cluster boundaries. As we face a rapidly emerging space of open-source projects that address the area of inter-cluster communication, the question arises as to how to assess and evaluate their impact on the application performance in edge environments.
In this work, we experimentally study and comparatively analyze for the first time three notable open source projects and tools for inter-cluster communication in edge networking and computing: Submariner, ClusterLink and Skupper. While there are several other tools for general inter-cluster communication, this study focuses on open-source solution that do not necessitate sidecar containers or a specific container network interface in the connected clusters, and as such allow for the highest degree of compatibility with existing infrastructure. To this end, we engineer an open-source testbed for inter-cluster communication tools. Its realization uses only production-grade open-source tools, while it provides the ability to vary the underlying network conditions to mimic an edge network infrastructure setting with various link qualities. The inter-cluster communication tools are implemented in a pluggable fashion, thus allowing for a comparative analysis. We also design a set of innovative case studies for novel distributed mobile applications, including use cases based on machine learning (ML), which necessitate inter-cluster communication. Finally, we provide measurements and evaluate the impact on application performance. In the measurements, we attempt to answer the following three research questions, which remain unaddressed by existing work for the scope of inter-cluster communication: 1. How does the distance between clusters impact the service response time of latency sensitive applications? 2. What is the combined effect of cluster distance and payload size on the application’s performance? 3. How do various inter-cluster communication tools comparatively perform for latency sensitive applications?
The remainder of this article is structured as follows: Section II provides the background on the inter-cluster communication tools and the analyzed use cases as well as the related work as state of the art. Section III presents the system architecture of the testbed used for performance benchmarking. Section IV describes the design and parameters of the tests executions. Section V discusses the results and Section VI concludes the paper.
II Background and Related Work
In this section, we first describe the related work on inter-cluster communication. We then briefly provide the background on Submariner, ClusterLink and Skupper, and finally provide assumptions on application case studies applied in the context of emerging mobile network distributed applications.
II-A Related work
Kubernetes is the de facto default tool for managing workloads in the cloud. It organizes computing resources in so-called clusters that can be formed based on different characteristics of nodes, administrative or management goals, or yet a combination of them. Dynamic clusters are formed based on aspects subject to continuous changes, like node proximity in scenarios with moving computing nodes [1]. Static clusters are usually based on hardware properties to form specialized groups, for example, to process tasks that demand hardware accelerators such as GPUs [4]. Such specialized clusters are employed to guarantee the compatibility of applications that demand specific resources since those devices might leverage enhancements to utilize specific resources like GPUs[5]. Additionally, scheduling and workload sharing can differ in those particular installations, which could lead to incompatibility with other Kubernetes clusters [6]. General clustering approaches are analyzed in [3], showing how dependent the selection objectives can be on the specific use case.
Distributed application and service function chains can be formed across multiple clusters. The applications are scheduled and scaled based on node mobility or network properties to minimize service latency[2, 7], increase cost-efficiency, reduce power consumption or boost robustness [3]. Multi-cluster deployments of applications are gaining attention because they address a variety of issues faced in distributed single-cluster environments. On the one hand, there are business reasons to rely on multiple cloud providers (i.e., , multi-cluster environments), such as preventing vendor lock-in and ensuring a global presence. On the other hand, there are technical reasons, such as the need for edge computing to fulfill stringent latency requirements of novel applications (e.g., extended reality) [8], creating scalability issues. In  [9], the authors show that the Kubernetes control plane can become a bottleneck when a certain cluster size is exceeded. This restriction limits the size of clusters since consensus has to be maintained across all member nodes. While the issue can be reduced by changing or improving the used key-value storage system, this does not provide a complete solution, and the problem can be expect to increase relative to the geographical spread of the nodes. Additionally, due to the inherent characteristics of edge nodes (i.e., constrained computing and storage capacity), some applications can only partially be executed in an edge cluster, which in turn requires communication with other clusters.
In general, research focusing on the inter-cluster communication problem to allow multi-cluster collaboration is work in progress. In [10], the authors perform an assessment of Submariner’s performance as well as the distributed application deployed on top of it with a focus on CPU utilization. However, this study lacks consideration of two critical dimensions: varying inter-node/inter-cluster latency and the amount of data being exchanged between the clusters. In [11] the authors explore inter-cluster communication by proposing a framework to deploy a 5G experimental network across distributed infrastructure formed by multiple geographically spread clusters. The authors evaluated the maximum achievable throughput in inter-cluster communications provided by Submariner using three different tunneling options (i.e., Wireguard, Libreswan, and VXLAN) and three different transport protocols (TCP, UDP and SCTP). These works focus on Submariner while our focus is on a comparative analysis in various scenarios with ClusterLink and Skupper.
Focusing on cloud-driven scenarios for mobile networks furthermore, paper [12] proposed a solution for multi-cluster management and connectivity employing two open-source tools: KubeFed and NSM. The Libreswan technology was used to tunnel traffic between clusters. The approach proposed utilized TCP, UDP, and SCTP as transport protocols and benchmarked basic data plane performance, collecting maximum throughput and latency metrics between containers. Since latency between nodes was static, real-world performance in the network edge would diverge significantly from these findings. A valuable extension of paper [12] would in fact be to compare the tool proposed with some of the production-grade tools that we use in our paper.
Related work in benchmarking of intra-cluster networking solutions, such as in [13] focuses on total performance within a single cloud cluster, which varies based on the hardware used. Paper [14] conducted an extensive survey of benchmarking performance at the edge and observed that further research is needed to measure the performance of edge networks efficiently. This survey reveals a gap in explicitly benchmarking orchestrators and other software that connects the infrastructure into an edge continuum. Indeed, many benchmarks are limited by physical aspects of edge devices (e.g., processing power, chipset throughput, storage, and memory). Our evaluation differs from previous works as, instead of absolute metrics like resource utilization, it focuses on the relative response time differences encountered through varying aspects of 1) underlying network, 2) payload size, and 3) direction of payload exchange between clusters.
II-B Inter-cluster Communication Tools
Tab. I lists inter-cluster communication tools that we implement in this paper; Submariner [^1] https://submariner.io/ (CNCF, RedHat) and Skupper [^2] https://skupper.io/, along with the emerging ClusterLink [^3] https://clusterlink.net/ (IBM) [15]. While all three of them tackle the same challenge that is of inter-cluster service communication, each tool has unique features that set it apart. Skupper, being a layer 7 interconnect, does not need any further configuration of the CIDR ranges that are used to assign IP addresses to pods or services, making it scale very well with comparably low configuration effort. Submariner on the other hand works on layer 3 and employs a broker for connection establishment, leading to only this one component having to be publicly accessible; it can potentially even be hosted on hardware that is not part of the communicating clusters. This allows for cluster-to-cluster communication between dynamic endpoints through NAT as well as firewalls. Using HTTP, ClusterLink shares the benefits of layer 7 connections but also provides authentication as well as granular permission and access management on workload, service and cluster level, allowing for a high degree of security and low levels of implicit trust between endpoints. All three solutions are written in Go, provide state-of-the-art encryption out of the box and are deployed on top of the respective Kubernetes clusters without the need for modifications of the individual node’s operating systems.
TABLE I: Comparison of cluster connectivity tools
Tool name Deployed Network Protocol Features
version driver
Skupper.io 1.7.3 mTLS AMQP simple configuration
Submariner 0.18.0 Libreswan IPsec NAT compatible
VXLAN VXLAN
Wireguard Wireguard
ClusterLink v0.4.0 mTLS HTTP ACLs
II-C Applications
The performance evaluation of the inter-cluster communication tools implemented uses key performance indicators (KPIs) and characteristics of the emerging mobile computing applications. We choose case studies of two classes of applications: industrial automation and cellular Vehicle-to-Everything (C-V2X). They are illustrative of the real-world verticals that require robust inter-cluster communication, i.e.,
II-C1 Remote Access, Monitoring, and Maintenance
In an industrial automation setting, numerous sensors installed on production machines distributed across a factory continuously collect data about the status of machines in the production line. A massive amount of raw data is cached in a cluster of edge nodes for preprocessing that aggregate the data where possible. The result of the local processing is sent to another cluster for evaluation (e.g., using machine learning), with the result potentially being used to instruct a control center to perform the required maintenance [16, 17, 20, 21].
II-C2 Vehicle Decision Assist
A specific vehicle that is part of a fleet of vehicles platooning on the road decides to overtake the vehicle in front, for example to optimize fuel consumption, avoid breaking when a preceding vehicle leaves the platoon at an upcoming exit, or because the vehicle in front is facing a mechanical problem. Due to road and traffic conditions, the vehicle cannot overtake easily, so the platoon asks for external information to assist with the maneuver. The exchange of information happens between the cluster formed by vehicles in the platoon and a cluster of Road Side Units or oncoming vehicles of the opposite direction (e.g., another platoon). The vehicles share speed, length, and position information to guarantee a safe overtaking. Since reliability and safety is of utmost importance, the overall system consumes information from multiple systems. In this use case, multi-cluster communication is required since the platooning vehicles per direction and the roadside units supporting the vehicles would all be grouped in different clusters [18, 19].
Table II illustrates the extracted measurable KPIs for the applications chosen, including references to the literature that we base these KPIs on.
TABLE II: Average payload size and max. response time per use case
Use case Message payload Max. response
size (byte) time (ms)
Remote Access, Monitoring, 60 - 1000000 1000
and Maintenance[16, 17, 20, 21]
Vehicle Decision Assist[18, 19] 1000 100
III System Architecture and Testbed Design
Fig. 1 illustrates our testbed. It includes four physical machines with the specifications listed in Table III, all of them are equipped with Gigabit interfaces and connected through a single Gigabit Ethernet switch. They represent off-the-shelf, but well-equipped consumer hardware which we assume to roughly represent the computational capacity of power-restricted (<=<=< = 100 Watts) edge nodes with small form factors that could be deployed in constrained spaces, e.g., in curb cabinets, cars or industrial machines.
Figure 1: Testbed design
TABLE III: Testbed Hardware
Cluster Role CPU Memory
cluster 1 headnode Intel Core i5-7600 @ 3.5 GHz 16 GB
worker Intel Core i5-10600 @ 3.3 GHz 16 GB
cluster 2 headnode Intel Core i5-7600 @ 3.5 GHz 32 GB
worker Intel Core i7-6700 @ 3.4 GHz 32 GB
The software stack used consists of hierarchical layers, as seen in Fig. 1. Since we publish our code base on Github [^4] https://github.com/michalke-it/inter-cluster-benchmarking, we design the testbed exclusively around open-source tools, akin to [22]. The functionalities provided by each layer are described below. This environment guarantees a common baseline for all three tools. Most notably, our testbed allows for the manipulation of network properties like delay, jitter, and packet loss through software, while keeping all other factors static, thus enabling comparative performance measurements.
III-A Operating System and Kubernetes
We deploy a linux-based operating system to provide basic functionality as well as all prerequisites needed by the container orchestrator. Here, we chose Debian bookworm due to its server-grade stability and wide acceptance, which makes it one of the most tested platforms for edge and cloud tools. We also employ the tool netem for manipulation of the Ethernet interface by artificially delaying or dropping frames, therefore emulating delay, jitter and packet loss as experienced relative to geographical distance or connection quality.
To provide a container orchestration layer, we choose the k3s Kubernetes distribution meant to be used for lightweight and edge devices, v1.30.3+k3s1. It provides the majority of features available in the regular Kubernetes version while having a much smaller footprint. Two clusters are formed with one worker and one headnode each and a total of four different IP ranges between the two clusters for their respective services and pods. As can be seen in Fig. 1, the worker nodes run the application and load generator respectively, while the headnodes are used exclusively to run each cluster’s control plane. This separation ensures that the impact of control plane tasks on the workers is minimized, while also replicating real-world use cases where at least one node per cluster has to be available via a defined address to allow for management and negotiation of connection endpoints.
III-B Inter-Cluster Communication
Two clusters are connected through the inter-cluster communication tool under test, namely Submariner, ClusterLink, and Skupper, as shown in the light blue box on Fig. 1. This setup allows services in cluster 2 to access exposed services in cluster 1. The inter-cluster communication tools are deployed as Kubernetes yaml definitions and are responsible for service discovery, service address resolution, and data transport across the connected clusters.
III-C Applications
The three example applications are functions served on top of the Knative [^5] https://knative.dev framework which handles requests and provides HTTP endpoints. Their source code is published on Github [^6] https://github.com/fcarp10/serverless-functions. While all of these applications are deployed at the same time, they are interacted with sequentially. Automatic scaling in Kubernetes is deactivated since we do not aim to measure total achievable performance on the given hardware but instead relative response time differences. The functionality of the respective functions can be summarized as follows:
III-C1 payload-echo
This function accepts any payload and echoes the same message back to the sender by setting the response body to that received payload. It is used in testing our Remote Maintenance (ML Inference) use case, since it emulates the transfer of aggregated data from one cluster to another with an equally sized response. In the test, the result shall be transmitted back to the sending cluster, which is the behavior that this function exhibits.
III-C2 payload-recv
This function receives any payload sent as the HTTP body and acknowledges it with an HTTP 200 response without body. It is used for the Remote Maintenance (ML Training) application, since we can see the same data transfer here as in the previous function with the difference that the data is not transferred back. This behavior would be expected whenever an ML model is trained, since the sending cluster would not have to wait for the training to complete.
III-C3 static-reply
This function simply replies to any requests with a static message, regardless of the request body. We use it to mimic the Vehicle Decision Assist application, where payload data is only received and not sent. Here, one cluster queries the information from another cluster without submitting any payload itself.
III-D Load Generator
To generate requests to the application, we utilize Jmeter in version 5.5, a load testing tool that allows for generating parallel and sequential HTTP requests. We deploy the tool as a container on top of Kubernetes to emulate the communication between two containerized services deployed on different clusters. The tool measures the time it takes between sending the request and receiving the reply and saves it to a csv file.
IV Evaluation Setup
The main challenge in the evaluation setup is to create the testing environment and applications as close as possible to the real world use case so that meaningful results for the application response times can be obtained. To this end, and based on the description provided in Section II, we map the analyzed communication patterns to our example applications, and design the matching Jmeter tests to evaluate them. We then gather the response times achieved through the different tools across a set of six different infrastructure scenarios that mimic edge environments with different characteristics. To create a baseline, we start with a test without any interface manipulation, leaving only the delay, jitter and packet loss experienced physically by the transmission. This establishes a reference value for each tool when running on our provided hardware and software stack.
We then start introducing delay, jitter and packet loss to the frames transmitted through the Ethernet interface, emulating a transmission over higher distances or lower quality connection media between all nodes as listed in Tab. IV. The naming of each parameter set is based on the additional delay introduced to an average round trip time between any two nodes; e.g. ,10ms in edg10. We base our values on the measurements described by the authors in [8] and extrapolate them to gain a set of test values.
Each Jmeter test is configured with base parameters like the number of sequential requests (40), the number of parallel threads (1) and the reuse of TCP sessions (enabled). Every test is executed across all six scenarios and repeated ten times on each of them to guarantee reproducible results. Before every run, the whole infrastructure is reverted to a clean, baseline state to prevent cross-contamination of the results.
TABLE IV: Emulated link parameters per scenario
Scenario Delay Jitter Packet Loss
lcl +0 ms +0 ms +0%
edg2.5 +1.25 ms +0.25 ms +0.02%
edg5 +2.5 ms +0.5 ms +0.04%
edg10 +5 ms +1 ms +0.08%
edg15 +7.5 ms +1.5 ms +0.12%
edg20 +10 ms +2 ms +0.16%
IV-A Applications
All tests are conducted based on Fig. 2, including the request and response direction. Tab. V details the payload size parameters of the files used to generate the requests/responses.
IV-A1 Remote maintenance inference
In this class of applications, the payload-echo function emulates the transmission of aggregated data to a specialized cluster for processing. The destination cluster communicates back to the aggregation cluster. Therefore, data has to be transmitted in both directions, where, depending on the specific data and actuators, the amount per direction could fluctuate significantly. For simplification, our destination cluster echoes back the received data, which should approximate the system behavior described in the use case. In Fig. 2, the payload size of the request and the response are therefore identical. During the execution of this test, the KPI for the response time is considered to be 1000 ms. Once this cannot be ensured by the different tools, potentially catastrophic failure in a production system can be assumed.
Figure 2: Data exchange across clusters
TABLE V: Payload size of data exchange between clusters
Use case Payload size
request response
Remote maintenance 60B, 1KB, 2KB, 10KB, 60B, 1KB, 2KB, 10KB,
inference 100KB, and 1MB 100KB, and 1MB
Remote maintenance 60B, 1KB, 2KB, 10KB, 0B
training 100KB, and 1MB
Vehicle decision 0B 1000B
assist
IV-A2 Remote Maintenance (ML Training)
In this test, using the function payload-recv, we emulate the behavior of the system sending data for training an ML model. The data is assumed to have been aggregated, resulting in variable-sized data in a cluster that is sent to a second cluster. Since the data is used for ML training, the destination cluster, cluster 1 in Fig. 2, only acknowledges receiving the data without any payload. The payload of the request ranges from 60B to 1MB, while the payload of the response is zero bytes (0B) as detailed in Table V. Since machine learning training has high tolerance regarding the response time, depending on the application, we choose an arbitrary maximum response time of 1000 ms. This allows us to compare the effect of a simple acknowledgement to the extensive application response of the previous test.
IV-A3 Vehicle decision assist
Here we use the static-reply function. A car in a platoon is overtaking and coordinating with either the roadside units of the section, or alternatively/complementary an oncoming platoon of vehicles. It has to receive data from that other cluster, e.g., platoon length, speed, and position, with low tolerance for delay. Therefore, we assume an empty payload for the request, which aims to trigger a response with a fixed payload size of 1000B, as shown in Table V.
As in all tests, we evaluate the results to determine with which inter-cluster network parameter set the response times break the maximum value to ensure save operation. Here, we set this value to 100 ms before the information is deemed so outdated that it loses its meaning and endangers safe operation of the machines and safe transportation of the passengers.
V Measurements and Results
In this section, we present the measured values and the results achieved through the tests described previously. Our measured application response times are defined as, i.e.,
$T_{total}=T_{reqSend}+T_{hostProc}+T_{appProc}+T_{replySend}$
where $T_{reqSend}$ is the request sending time including possible retransmissions, $T_{hostProc}$ is the host processing time at the destination cluster which includes the time taken by (de-)fragmentation and reassembly of the request. $T_{appProc}$ is the delay introduced by the processing of the application including potential processing of the payload, which depends on the parameter size in bytes. In practice, the differences introduced by the processing of different parameter sizes in our experiments are compensated for by the baseline measurements (loc0 scenario) in our results. Finally, $T_{replySend}$ is the reply sending time including possible delays introduced by retransmission mechanisms in the network layer.
During the evaluation, we focus especially on the two sending times since they are specific to the inter-cluster communication tools and point out differences between them per tool.
V-A Remote Maintenance (ML Inference)
Figure 3: Results for Remote Maintenance (Inference) 1MB
Fig. 3 shows the response times achieved via the respective inter-cluster communication tools for the ML inference in the Remote Maintenance use case with a payload of 1MB, the largest one in our value set. In this case study, the data sent is echoed back to the sender, generating a response of the same size as the request. We can observe a continuous, exponential increase of response times relative to the deteriorating networking conditions when using Submariner and Skupper, with Skupper achieving roughly two thirds of the response time of Submariner in our worst case. While Clusterlink is subject to the same trend, it seems to better deal with the additional delay imposed on the network connection between clusters, showing response times almost 50% lower than Skupper when using our lowest-quality links. In general, all tools were able to handle the request and response without their median response times exceeding the 1000 ms limit for the use case as defined in Table II. Submariner however occasionally exceeds the tolerable upper bound slightly when the parameter set with 20 ms of delay is introduced to the connection, which indicates that the tool is more sensible to link delay, delay variance and packet loss when compared to other tools.
Figure 4: Results for Remote Maintenance (Inference) 100KB
For smaller payloads, ClusterLink and Submariner performed similarly until the eg10 value set, with Submariner being slightly outperformed by ClusterLink in most cases as depicted in Fig. 4. However, the difference is subtle and in consideration of run to run variance, their performance can be considered equal here. Skupper, while initially delivering comparable performance to the other tools breaks away in the edg15 scenario by delivering ever higher response times, by almost 20% in the edg20 scenario, indicating a downside of using AMQP in this environment. From a response time perspective, this makes Skupper the last choice out of the three tools for inter-cluster communication when dealing with applications that need to send medium-sized payloads to a service that in turn replies with small payloads.
Figure 5: Results for Remote Maintenance (inference) edg15
In Fig. 5, we analyze the payload size’s impact on the tools’ performance for the remote maintenance inference use case in the scenario with the worst conditions that still allow all tools to perform below the boundary; edg15. We can observe similar performance for the three tools, with mean values below 100 ms for payloads up to 100 KB. With the payload size increased to 1MB, we see how the performance of the tools diverges, with drastic increases of the response times of Submariner, Skupper scaling slightly better, while ClusterLink’s jump in response times is significantly less pronounced.
We can therefore conclude that for workloads where small amounts of data are sent to remote services that reply with a similar amounts of data, the three tools perform similarly to each other. For medium and larger data, i.e. 100 KB and more, the tool selection becomes a decision of significant impact.
V-B Remote Maintenance (ML Training)
Fig. 6 shows the results of the ML training case studies for 1MB payload size, which, differently from the inference use case, does not consider communication back to the request source besides acknowledgments. Again, we can observe ClusterLink and Skupper behaving similarly to each other with slightly lower medians for ClusterLink, while it now also exhibits a slightly higher spread. Similarly to the ML inference for the remote maintenance use case with payloads of 1 MB, Submariner performs worst among the tested tools. Here, it presents about four times the median values of ClusterLink and Skupper in the edg20 scenario along with a high standard deviation in the observed data. In contrast to the previous use case, Submariner exceeds the boundary of 1000 ms already in the edg10 scenario with some measurements and in the edg15 scenario with its median value. ClusterLink and Skupper however remain below the threshold with their median values but slightly exceed the boundary in the edg20 scenario with a small portion of the observed samples.
Figure 6: Results for Remote Maintenance (Training) 1MB
The results for experiments with smaller payloads of 60 Bytes are presented in Fig. 7. The plots show almost identical performance for the three inter-cluster communication tools, to the level of some even reaching identical median values in the edg15 and edg20 scenario.
Figure 7: Results for Remote Maintenance (Training) 60B
In Fig. 8, we show the impact of varying payload sizes on the tools’ performance for the remote maintenance training use case under the worst networking conditions that still allow all tools to achieve median response times below the boundaries. We can observe similar performance for all three tools until the payload size exceeds 2 KB. Higher payloads lead to Submariner’s performance degrading significantly quicker than for ClusterLink and Skupper in relation to the payload size, indicating a clear advantage of the mTLS-based implementations over the IPsec-based Submariner due to less overhead of this approach. ClusterLink then brings this overhead down further by utilizing HTTP instead of AMQP for transport, resulting in even better performance than Skupper.
Curiously, the worst performance of Submariner that we observed was when data was not sent back to the tester container (or client), apart from acknowledgments. We can therefore conclude, that Submariner’s implementation seems to have suboptimal handling of HTTP POST responses when it comes to larger payloads, which makes it unsuitable for this application when network latency is above 7.5 ms in each direction (assuming a symmetric link, therefore 15 ms for round trip).
Figure 8: Results for Remote Maintenance (Training) edg15
V-C Vehicle Decision Assist
Fig. 9 shows the results for the vehicle decision assist class of applications. ClusterLink and Skupper again present very similar performance with their median values being identical across almost all scenarios. Submariner however seems to consistently perform slightly better (1 ms lower median response times) than the other tools. When handling these small payloads, according to our results, the inter-node connection quality only has a linear influence on the total response time. The response times of all three tools remain inside the set boundary of 100 ms, making each of them a suitable choice for this class of application. A further increase of the distance between the nodes beyond a round trip time of 20 ms would be unrealistic, since such a high latency for devices that are in danger of physically touching each other is highly unlikely.
Figure 9: Results for Vehicle Decision Assist
V-D Discussion
Since this is an experimental research work, the applicability of our results is limited in the following ways:
V-D1 Hardware
Our measurements strongly depend on the individual hardware, and we cannot completely exclude the possibility that the three tools would score in a different order on different hardware due to better encoding or encryption performance. We however believe, that the tested hardware represents a reasonably broad group of edge devices and therefore should be able to provide a general idea of which tools suit which circumstances.
V-D2 Protocols
The results might not apply to streaming protocols used by real-time applications. With their way of data partitioning and transmission, they should be evaluated separately.
V-D3 TCP Connection Reuse
Our tests have had the “reuse TCP connections” activated in JMeter which we believe is a reasonable approach for traffic between two clusters. This however does not represent classes of applications where multiple clusters or services communicate with a common destination cluster. Individual results for such circumstances can therefore differ from what we have presented here.
V-D4 Payload Size
The step increase of our payloads is quite large to allow for displaying a wide variety of different values. This might or might not reflect the performance to be expected with other applications and therefore should be repeated with values closer to the intended data sizes.
V-D5 Tool Maturity
At the time of testing, Clusterlink is available in v0.4.0, meaning it has not yet hit a stable release target. Through code improvements, its performance might be subject to change, while Submariner and Skupper are established and production-ready tools.
V-D6 Inter-node and Inter-cluster Latency
In this work, we assume equal distance between two nodes of the same cluster and between different clusters. In reality, the inter-cluster distance would often be higher than the distance within clusters, depending on the metric that the cluster is formed on.
V-D7 Submariner Network Driver
We only explored Submariner’s libreswan (IPsec) driver, while other research [11] has shown that its performance could vary when different drivers are used.
VI Conclusion
In this work, we evaluated for the first time the impact of inter-cluster communication on edge computing performance by using three prominent, open source inter-cluster communication projects and tools, i.e., Submariner, ClusterLink and Skupper. We developed an open source testbed that integrates these tools in a modular fashion, and experimentally benchmarked the sample applications, including the ML class of applications, on their performance under varying networking conditions.
In answering the research questions stated upfront, we proved that the response times have a linear relation to the geographical distance between all nodes when the transmitted data is low in size, while it can approach exponential growth with large payload sizes. We also observed that the combined impact of geographical spreading of nodes and increasing payload size on the considered applications leads to exponential growth of application response times, at least for data being transferred into both directions and for data being sent. For the three tools analyzed, we found that ClusterLink outperformed Submariner and Skupper whenever larger data was exchanged with a service, Submariner performed slightly better for small payloads that are received via HTTP GET requests and Skupper performed decently unless medium-sized payloads had to be pushed via HTTP POST requests.
In more general terms, we can say that for inter-cluster communication, creating network layer tunnels like Submariner seems to provide higher efficiency for small payloads but suffers when dealing with large amounts of data as soon as significant delay, variance and packet loss is experienced on the connecting link. Solutions with the ability of proxying TCP connections are at advantage here, delivering lower overall response times. Furthermore, use cases where data has to be sent to another service with significantly sized replies seem to suffer when using Skupper or Submariner in edge networks with suboptimal inter-node connections.
It should be noted that the individual choice of software for a real-world setting should however always evaluate not just the performance, but also the feature set of the tools. For instance, if strong security is needed, ClusterLink’s authentication and access control might prevail, just as Submariner’s flexibility for connection establishment might make it more a more attractive choice. Also, when simplicity of the configuration setup is the main concern, Skupper might be chosen as the simplest solution.
Acknowledgment
This work was partially supported by the German Federal Ministry of Education and Research (BMBF) project 6G-ANNA, grant agreement number 16KISK100, and the project EU HORIZON ICOS under GA no. 101070177.